• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name T.Chang

    Log entry time 16:05:14 on June28,2002

    Entry number 86966

    keyword=shift summary

    8:00 Hall-A in controlled access. Change beam energy to 3-pass.
    10:30 Finished target survey. Move the left spectrometer to 37 degree and the right spectrometer to 36.8 degree. Trigger downloaded.
    11:15 Beam permit. MCC says there are problems with BCM. Need about 30 mins
    for the temperature to stablize.
    11:20 Move to 9-foil carbon target. Turn off the raster.
    11:35 Beam back at 35 uA. run 1880 C-fiol optics run w/o raster.
    11:50 Asked for raster. start run 1881.
    12:10 Recycle Q2 and Q3 on both arms.
    12:25 Asked for bigger raster(1mmx2mm). start run 1882(optics target).
    12:40 start run 1883, LH2 production run in config A.
    13:05 start run 1884, LH2 production run in config A.
    13:30 done with config A. Switch to config B.
    13:50 start run 1885, LH2 production run in config B. stopped due to mcc wants to do some BCM calibration.
    14:15 start run 1886 in config B. end run at 14:35.
    14:35 start run 1887 in config B. end run at 14:55.
    14:55 start run 1888 in config B. end run at 15:05.
    15:05 done with config B. Switch to config C.
    15:30 start run 1889 in config C. end run at 15:50.
    15:50 start run 1890 in config C. end run at 16:05.